OpenStudy (anonymous):

A sample of gas in a balloon has an initial temperature of 11 degree Celsius and a volume of 1490L. If the temperature changes to 86 degree Celsius, and there is no change of pressure or amount of gas, what is the new volume,V2, of the gas?

OpenStudy (rogue):

Remember to convert temperature to kelvin. \[\frac {V_1}{T_1} = \frac {V_2}{T_2}\]

OpenStudy (anonymous):

\[volume \alpha temperature\] \[V \times T = K\] Find the constant...first to make it easy convert the Celcsius degrees to kelvin then substitue the temperature and the constant to get the volume